Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold for $860. Selected data for the company’s operations last year follow:

Units in beginning inventory 0
Units produced 320
Units sold 285
Units in ending inventory 35
Variable costs per unit:
Direct materials $ 135
Direct labor $ 355
Variable manufacturing overhead $ 30
Variable selling and administrative $ 15
Fixed costs:
Fixed manufacturing overhead $ 64,000
Fixed selling and administrative $ 27,000

The absorption costing income statement prepared by the company’s accountant for last year appears below:

Sales $ 245,100
Cost of goods sold 205,200
Gross margin 39,900
Selling and administrative expense 31,275
Net operating income $ 8,625

Required:

1. Under absorption costing, how much fixed manufacturing overhead cost is included in the company's inventory at the end of last year?

2. Prepare an income statement for last year using variable costing.
Answer #1
1
Fixed manufacturing overhead cost included in inventory 7000 =64000/320*35
2
Variable costing Income Statement
Sales 245100
Variable expenses:
Variable cost of goods sold 148200 =285*(135+355+30)
Variable selling and administrative expenses 4275 =285*15
152475
Contribution margin 92625
Fixed expenses:
Fixed manufacturing overhead 64000
Fixed selling and administrative expenses 27000
91000
Net operating income 1625